# SlideAgent: Hierarchical Agentic Framework for Multi-Page Visual Document Understanding

### v2

@@ Abstract
Multi-page visual documents such as manuals, brochures, presentations, and posters convey key information through layout, colors, icons, and cross-slide references. While large language models (LLMs) offer opportunities in document understanding, current systems struggle with complex, multi-page visual documents, particularly in fine-grained reasoning over elements and pages. We introduce SlideAgent, a versatile agentic framework for understanding multi-modal, multi-page, and multi-layout documents, especially slide decks. SlideAgent employs specialized agents and decomposes reasoning into three specialized levels–global, page, and element–to construct a structured, query-agnostic representation that captures both overarching themes and detailed visual or textual cues. During inference, SlideAgent selectively activates specialized agents for multi-level reasoning and integrates their outputs into coherent, context-aware answers. Extensive experiments show that SlideAgent achieves significant improvement over both proprietary (+7.9) and open-source models (+9.8).
@@ 1 Introduction
Visual documents–from earnings reports and academic lectures to business strategy presentations–are ubiquitous, conveying ideas not only from text, but also from the intricate interplay of layout, icons, visual hierarchy, and cross-page relationships. These documents are central to high-stakes domains such as finance, science, and technology. Accurately interpreting them thus remains a pressing challenge.
Figure 1: When given the full page, the LLM miscounts the number of product mix categories. After isolating the chart, it correctly identifies all eight categories, highlighting the importance of accurate element parsing.
@@
Accurately interpreting these multi-page, multi-modal artifacts remains challenging. Recent advances in multimodal large language models (MLLMs) have accelerated document understanding (Verma et al., 2024), yet three gaps remain: 1) Scalable Fine-Grained Reasoning. State-of-the-art MLLMs process a limited number of images at once (Liu et al., 2023) and tend to treat each page holistically, missing fine-grained, element-level cues required for user queries (Faysse et al., 2024; Tanaka et al., 2025). In Figure 1, an LLM can miscount chart segments on a cluttered page of a slide deck, but correctly identifies all segments once the relevant chart is cropped–highlighting the importance of element-level parsing for latent reasoning abilities. 2) Domain-Specific Visual Semantics. Most MLLMs are pre-trained on natural images (Wu et al., 2024), lacking exposure to domain-specific diagrams, financial charts, or scientific plots. Consequently, they struggle with the specialized language of visual documents (Cho et al., 2024). For example, logos appear on every page, reinforcing brand identities but do not offer additional content. Color schemes encode categorical information (red for losses and green for gains in financial reports). Icons convey abstract concepts (lightbulbs for innovation, arrows for causal relationships); and spatial positioning signals importance (centered elements typically matter more than corner annotations). falter in spatial reasoning Wang et al. (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al. (2025); Polak and Morgan (2025). Low-resolution visual encoders in MLLMs (Liu et al., 2023) further miss details such as footnotes or superscripts. 3) Metadata-Free Integration. Many systems  Singer-Vine (2025); huridocs (2025); Rausch et al. (2021) rely on clean metadata–figure locations, hierarchy tags, embedded text layers–that can be unavailable or corrupted in real-world PDF. Users may take screenshots of PDF documents, scan copies of physical documents, export slides or documents as flattened PDFs, upload or share PDFs generated from software that strips out or does not preserve document structure. Recent metadata-free methods (Yu et al., 2025; Tanaka et al., 2025) address these by parsing only visual images without relying on the metadata, despite a performance gap.
@@
We present SlideAgent, an LLM-based agentic framework for fine-grained understanding of multi-page, varying-size visual documents. Inspired by the human information processing model (Lang, 2000; Naysmith et al., 2021), SlideAgent employs a hierarchical architecture with specialized agents at three levels: global (document-wide topics), page (page-specific features and cross-page relations), and element (fine-grained components such as charts, figures, and text blocks). During knowledge construction stage, SlideAgent parses layout and generates query-agnostic knowledge at each level. At inference stage, SlideAgent retrieves query-specific knowledge, enabling scalable fine-grained reasoning over relevant pages and elements. We benchmark SlideAgent and baseline models, demonstrating significant performance on both open-source and proprietary models. SlideAgent surpasses its LLM counterpart by 7.9 (proprietary) and 9.8 (open-source), respectively. We also show that SlideAgent enhances spatial reasoning, visual counting, and cross-element understanding, with results that are highly interpretable.
@@ 2 Method
Figure 2: SlideAgent generates knowledge about input slide decks in a hierarchical manner at 3 levels: global, page, and element. At each level, specialized agents generate query-agnostic knowledge during knowledge construction, then retrieve and reason over query-specific textual & visual knowledge during inference stage. Sample knowledge $\mathcal{K}$ generated by SlideAgent is in Appendix Figure 7,8,9 and answers generated by the agents are in Figure 5.
@@ 2.1 Knowledge Construction Stage
The global agent $\mathcal{M}_{g}$ generates document-level knowledge $\mathcal{K}_{g}=\mathcal{M}_{g}(\mathcal{P})$, capturing the overall summary, objectives, and narrative flow of the document. This layer establishes overarching themes to support high-level reasoning about the document’s purpose. Since visual documents are often large and LLMs have a limited capacity for processing visuals, SlideAgent samples the first three pages to generate $\mathcal{K}_{g}$ (example in Appendix Figure 7).
@@
For each page $p_{i}\in\mathcal{P}$, the page agent generates page-level knowledge $\mathcal{K}_{p}$ in a sequential manner, conditioned on the page’s visual content $v_{i}$, the global knowledge $\mathcal{K}_{g}$, and the knowledge from the preceding page $\mathcal{K}_{p}^{i-1}$:
$\mathcal{K}_{p}^{i}=\mathcal{M}_{p}(v_{i},\mathcal{K}_{g},\mathcal{K}_{p}^{i-1}),i\in[1,|\mathcal{P}|].\quad\mathcal{K}_{p}^{0}=\emptyset$ (1)
The complete page-level knowledge $\mathcal{K}_{p}=\bigcup_{i=1}^{|\mathcal{P}|}\mathcal{K}_{p}^{i}$ provides an intermediate representation that captures page-specific content while linking them to the global context (sample in Figure 8). To ensure comprehensive understanding of all pages, $\mathcal{K}_{p}$ is subsequently used to refine $\mathcal{K}_{g}$.
@@
LLMs often struggle with spatial reasoning over visual documents Wang et al. (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al. (2025); Polak and Morgan (2025). To address this, our element agent integrates external tools to explicitly capture the spatial and structural information of each page.
At the finest granularity, the element agent decomposes each page $p_{i}$ into a set of elements using a layout parsing pipeline $f:v_{i}\rightarrow\{(i,e_{j},b_{j},t_{j})\}_{j=1}^{M_{i}}$, where each element is represented by its page index $i$, verbatim text $e_{j}$, bounding box coordinates $b_{j}$, and element type $t_{j}$. $f$ integrates text detection, layout detection, and element classification, followed by post-processing to merge fragmented elements (Appendix A.1).
@@ 2.2 Inference Stage
Different queries require different perspectives to answer effectively. For instance, queries about global understanding, like “What is the overall theme of this presentation?” requires a broad overview from a macroscopic perspective and activates only the global agent. In contrast, a fact-based query, like “What is the revenue on slide 3?” requires detailed, slide-specific information and requires both the page and element agents. Leveraging too many agents may increase computation or introduce noise. Thus, the agent orchestrator first attempts to classify each query $q$ into one of 4 predefined categories, such as global understanding, fact-based direct queries, multi-hop reasoning, and layout & visual relationships. Each type corresponds to a question-answering strategy and a targeted set of agents (Appendix Table 7). For instance, global understanding queries activate only the global agent, as they require a broad overview, while fact-based queries trigger both the page and element agents to retrieve detailed, slide-specific information. If none of the predefined categories apply, the agent defaults to the “unknown” category, which activates all agents.
@@ 3 Experiments
Gemini 2.0 75.0 71.3 79.8 50.4 67.6 41.6 70.8 70.6 77.8
Gemini 2.5 83.8 78.3 91.8 51.1 71.4 41.2 76.2 75.8 100.0
Gemini 2.5-lite 71.2 60.8 87.0 47.3 58.1 41.9 57.0 56.6 68.3
Claude 4.1 78.4 74.3 82.3 61.0 81.4 52.3 56.5 54.8 73.3
Claude 3.5 62.5 68.3 54.6 52.5 80.2 39.5 48.5 49.5 29.6
@@
Multimodal RAG (Type 2) and Agentic Methods (Type 3)
@@
Table 1: Performance comparison of proprietary models on SlideVQA, TechSlides, and FinSlides. All baseline methods use GPT-4o for answer generation. Improvements over the base model GPT-4o is shown in green. The best and second best performance are highlighted in bold and underlined. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model (GPT-4o), and even outperforms stronger raw LLMs (e.g. Gemini-2.5).
@@
We benchmark SlideAgent against 3 types of baselines: 1) Multimodal LLMs: 15 LLMs from 8 model families, including proprietary models (GPT-4o (OpenAI, 2025), Gemini (Anil et al., 2023), Claude (Anthropic, 2025)) and open-source models (Llama-3.2 (Grattafiori et al., 2024), InternVL3 (Chen et al., 2024), Phi-3 (Abdin et al., 2024), Qwen2.5-VL (Bai et al., 2025)); 2) Multimodal RAG Methods: VisRAG Yu et al. (2025), VDocRAG Tanaka et al. (2025), and COLPALI Faysse et al. (2024); 3) Multi-agent Systems: ViDoRAG Wang et al. (2025a). We evaluate SlideAgent using two backbone LLMs, including both proprietary models (GPT-4o (OpenAI, 2025)) and open-source models (InternVL3-8B (Chen et al., 2024)). These models are chosen for their widespread use in state-of-the-art QA systems Yu et al. (2025); Jin et al. (2025); Cho et al. (2024). The parameter sizes, knowledge cutoff dates, and release dates are in Appendix Table 8. We use the text-based retriever SFR (Meng et al., 2024) due to its strong efficiency-performance Cheng et al. (2024); Jin et al. (2025). For models restricted to single-image input (e.g. LLaVA-v1.5 (Liu et al., 2023)), we concatenate the top 3 retrieved images into one following previous work Yu et al. (2025). In settings where ground-truth pages are available, we provide them as input to all models. Otherwise, each model receives as many retrieved pages up to its input capacity.
@@
Raw LLMs
Gemini 2.0 72.3 66.1 81.3
Gemini 2.5 86.9 81.1 95.4
Gemini 2.5 lite 71.7 62.6 87.0
Claude 4.1 36.3 35.4 38.9
Claude 3.5 31.5 30.8 37.0
@@
Table 2: Performance comparison among Gemini, Claude, GPT-4o, ViDoRAG, and SlideAgent models.
@@ 3.1 End-to-End Performance
Tables 1/2/3 and Appendix Table 5 demonstrate the end-to-end performance of SlideAgent across proprietary and open-source models.
@@
Llama 3.2 11B 42.9 43.3 42.3 41.4 52.5 36.2 23.3 23.2 26.2
Phi3 72.3 61.8 90.6 59.4 60.0 59.1 48.8 48.5 64.3
Qwen2.5 7B 79.5 70.5 94.3 59.3 52.5 65.9 53.6 52.5 85.7
Qwen2.5 32B 79.2 71.1 92.2 67.5 87.5 60.6 57.4 56.6 87.5
LLaVA 1.5 7B 36.8 22.4 79.0 23.3 12.5 38.7 10.7 10.1 16.6
LLaVA 1.5 13B 44.9 25.1 81.8 28.1 17.5 45.0 20.6 16.5 36.7
LLaVA 1.6 7B 50.9 37.3 82.6 34.4 37.5 32.2 12.2 12.1 17.8
LLaVA 1.6 13B 16.7 10.2 81.5 45.2 40.0 49.1 32.0 31.3 64.3
InternVL3 8B 63.0 56.5 74.1 55.4 57.5 54.4 49.8 49.5 64.3
Multimodal RAG (Type 2) and Agentic Method (Type 3)
@@
Table 3: Performance comparison of various models on SlideVQA, TechSlides, and FinSlides datasets. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model (GPT-4o).
@@
For both proprietary and open-source models, SlideAgent consistently outperforms all baseline methods (Type 2/3) and the base models across all metrics. For proprietary models (Table 1), SlideAgent improves overall accuracy by +7.9, numeric reasoning by +8.3, and lexical overlap by +6.5 on SlideVQA. Similar trends are observed for TechSlides (+7.5 overall; +12.3 F1) and FinSlides (+5.5 overall; +17.5 F1), indicating robust performance across diverse domains. For open-source models, the advantage remains pronounced: SlideAgent outperforms InternVL3-8B by +9.8 overall (+11.7 numeric), showing that our hierarchical, multi-agent design generalizes well across LLMs.
@@
SlideAgent consistently achieves the best or second-best performance among proprietary models (Tables 1 and  11). Notably, although the base model slightly lags behind stronger LLMs such as Gemini 2.5 in raw capability (e.g. 83.8 for Gemini 2.5 vs. 77.0 for GPT-4o in Table 1), SlideAgent’s structured reasoning pipeline fills this gap (84.9 overall for SlideAgent). For open-source models, SlideAgent achieves better performance across all models except for the Qwen2.5 family. While strong base models such as Gemini-2.5 and Qwen2.5 already exhibit advanced multimodal comprehension, SlideAgent is model-agnostic in nature and can be directly applied to these models to further enhance performance.
@@
Figure 4 presents the performance breakdown across query types, as defined in Table 7. SlideAgent improves question-answering across diverse query types, especially in multi-hop reasoning and visual/layout questions. The greatest improvement occurs on case 3 (multi-hop reasoning), with a 9.8 point improvement (67.4 to 77.2). This means explicitly guiding the model’s reasoning using generated knowledge $(\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e})$ significantly improves reasoning capabilities. A notable 7.7-point improvement (66.7 $\rightarrow$ 74.4) is also observed in visual/layout reasoning, demonstrating the benefit of fine-grained element-level reasoning and retrieval that multimodal LLMs hardly achieve. Both SlideAgent and its counterpart perform well on Case 2 (Fact-based Direct Queries), with only a modest 2.1-point improvement, reflecting the model’s ability to handle page-level reasoning with little space for further enhancement.

### v3

@@ Abstract
Multi-page visual documents such as manuals, brochures, presentations, and posters convey key information through layout, colors, icons, and cross-slide references. While multimodal large language models (MLLMs) offer opportunities in document understanding, current systems struggle with complex, multi-page visual documents, particularly in fine-grained reasoning over elements and pages. We introduce SlideAgent, a versatile agentic framework for understanding multi-modal, multi-page, and multi-layout documents, especially slide decks. SlideAgent employs specialized agents and decomposes reasoning into three specialized levels–global, page, and element–to construct a structured, query-agnostic representation that captures both overarching themes and detailed visual or textual cues. During inference, SlideAgent selectively activates specialized agents for multi-level reasoning and integrates their outputs into coherent, context-aware answers. Extensive experiments show that SlideAgent significantly improves accuracy over both proprietary (+7.9%) and open-source models (+9.8%).
@@ 1 Introduction
Visual documents–from earnings reports and academic lectures to business strategy presentations–are ubiquitous, conveying ideas not only from text, but also from the intricate interplay of layout, icons, visual hierarchy, and cross-page relationships. These documents are central to high-stakes domains such as finance, science, and technology, and small misunderstandings can cascade into costly decisions: analysts extract key metrics from earnings decks, engineers review tech specs to avoid implementation errors, and educators generate study materials whose inaccuracies can propagate at scale. Accurately interpreting them thus remains a pressing challenge.
Figure 1: Accurate element parsing is important for fine-grained visual document understanding. The MLLM miscounts product-mix categories when reasoning over the entire slide, but succeeds after isolating the chart.
@@
Accurately interpreting these multi-page, multi-modal artifacts remains challenging. Recent advances in multimodal large language models (MLLMs) have accelerated document understanding (Verma et al., 2024), yet three gaps remain: 1) Scalable Fine-Grained Reasoning. State-of-the-art MLLMs process a limited number of images at once (Liu et al., 2023) and tend to treat each page holistically, missing fine-grained, element-level cues required for user queries (Faysse et al., 2024; Tanaka et al., 2025). In Figure 1, an MLLM can miscount chart segments on a cluttered page of a slide deck, but correctly identifies all segments once the relevant chart is cropped–highlighting the importance of element-level parsing for latent reasoning abilities. 2) Domain-Specific Visual Semantics. Most MLLMs are pre-trained on natural images (Wu et al., 2024), lacking exposure to domain-specific diagrams, financial charts, or scientific plots. Consequently, they struggle with the specialized language of visual documents (Cho et al., 2024). For example, logos appear on every page, reinforcing brand identities but do not offer additional content. Color schemes encode categorical information (red for losses and green for gains in financial reports). Icons convey abstract concepts (lightbulbs for innovation, arrows for causal relationships); and spatial positioning signals importance (centered elements typically matter more than corner annotations). However, current MLLMs falter in spatial reasoning Wang et al. (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al. (2025); Polak and Morgan (2025). Low-resolution visual encoders in MLLMs (Liu et al., 2023) further miss details such as footnotes or superscripts. 3) Metadata-Free Integration. Many systems  Singer-Vine (2025); huridocs (2025); Rausch et al. (2021) rely on clean metadata–figure locations, hierarchy tags, embedded text layers–that can be unavailable or corrupted in real-world PDF. Users may take screenshots of PDF documents, scan copies of physical documents, export slides or documents as flattened PDFs, upload or share PDFs generated from software that strips out or does not preserve document structure. Recent metadata-free methods (Yu et al., 2025; Tanaka et al., 2025) address these by parsing only visual images without relying on the metadata, despite a performance gap.
@@
We present SlideAgent, an MLLM-based agentic framework for fine-grained understanding of multi-page, varying-size visual documents. Inspired by the human information processing model (Lang, 2000; Naysmith et al., 2021), SlideAgent employs a hierarchical architecture with specialized agents at three levels: global (document-wide topics), page (page-specific features and cross-page relations), and element (fine-grained components such as charts, figures, and text blocks). During knowledge construction stage, SlideAgent parses layout and generates query-agnostic knowledge at each level. At inference stage, SlideAgent retrieves query-specific knowledge, enabling scalable fine-grained reasoning over relevant pages and elements. We benchmark SlideAgent and baseline models, demonstrating significant performance on both open-source and proprietary models. SlideAgent surpasses its MLLM counterpart by 7.9% (proprietary) and 9.8% (open-source) in accuracy, respectively. We also show that SlideAgent enhances spatial reasoning, visual counting, and cross-element understanding, with results that are highly interpretable.
@@ 2 Method
Figure 2: SlideAgent generates knowledge about input slide decks in a hierarchical manner at 3 levels: global, page, and element. At each level, specialized agents generate query-agnostic knowledge during knowledge construction, then retrieve and reason over query-specific textual & visual knowledge during inference stage. Sample knowledge $\mathcal{K}$ generated by SlideAgent is in Appendix Figure 10,11,12 and answers generated by the agents are in Figure 8.
@@ 2.1 Knowledge Construction Stage
The global agent $\mathcal{M}_{g}$ generates initial document-level knowledge $\mathcal{K}_{g}^{(0)}$, capturing the overall summary, objectives, and narrative flow of the document. This layer establishes overarching themes to support high-level reasoning about the document’s purpose. Since visual documents are often large and MLLMs have a limited capacity for processing visuals, SlideAgent samples the first three pages to generate $\mathcal{K}_{g}^{(0)}$ (Appendix Figure 10).
@@
For each page $p_{i}\in\mathcal{P}$, the page agent generates page-level knowledge $\mathcal{K}_{p}$ in a sequential manner, conditioned on the page’s visual content $v_{i}$, the initial global knowledge $\mathcal{K}_{g}^{(0)}$, and the knowledge from the preceding page $\mathcal{K}_{p}^{i-1}$:
$\mathcal{K}_{p}^{i}=\mathcal{M}_{p}(v_{i},\mathcal{K}_{g}^{(0)},\mathcal{K}_{p}^{i-1}),i\in[1,|\mathcal{P}|].\mathcal{K}_{p}^{0}=\emptyset$ (1)
The complete page-level knowledge $\mathcal{K}_{p}=\bigcup_{i=1}^{|\mathcal{P}|}\mathcal{K}_{p}^{i}$ provides an intermediate representation that captures page-specific content while linking them to the global context (sample in Figure 11). To ensure comprehensive understanding of all pages, $\mathcal{K}_{p}$ is subsequently used to refine $\mathcal{K}_{g}^{(0)}$ through a single-pass fieldwise rewrite:
$\mathcal{K}_{g}=\mathcal{M}_{g}^{\mathrm{refine}}(\mathcal{K}_{g}^{(0)},\mathrm{concat}(\mathcal{K}_{p}^{1},\ldots,\mathcal{K}_{p}^{|\mathcal{P}|})).$ (2)
The global agent regenerates each global field from the complete page-level evidence, using $\mathcal{K}_{g}^{(0)}$ only as a weak prior for terminology and early context. This rewrite can overwrite earlier hypotheses when later pages provide stronger evidence, reducing bias toward the first pages. The refinement prompt is shown in Appendix Table 13.
@@
MLLMs often struggle with spatial reasoning over visual documents Wang et al. (2024b, a), failing to locate visual elements Sharma and Vats (2025); Bhattacharyya et al. (2025); Polak and Morgan (2025). To address this, our element agent integrates external tools to explicitly capture the spatial and structural information of each page.
Case Example Agents
Global Understanding Asks about the overall theme, purpose, or general summary of the entire presentation. "What is the main topic of the presentation?", "What is this deck about?" Global
Fact-based Direct Query. Asks for specific facts, data, or information from particular slides. "What is the revenue reported on slide 7?", "Which slide shows the product roadmap?" Page, Element
Multi-hop Reasoning Requires comparing information across multiple slides or elements. "Compare revenues in slide 5 and slide 10", "How do Sweden and Denmark compare?" Global, Page, Element
Layout / Visual Relationship Asks about visual relationships, positioning, or layout elements. "What does the diagram below the table on slide 12 illustrate?", "Is the color red used to denote negative performance?" Element
Uncertain If the query is unclear, use all agents to answer the query. \ Global, Page, Element
At the finest granularity, the element agent decomposes each page $p_{i}$ into a set of elements using a layout parsing pipeline $f:v_{i}\rightarrow\{(i,e_{j},b_{j},t_{j})\}_{j=1}^{M_{i}}$, where each element is represented by its page index $i$, verbatim text $e_{j}$, bounding box coordinates $b_{j}$, and element type $t_{j}$. $f$ integrates text detection, layout detection, and element classification, followed by post-processing to merge fragmented elements (Appendix B.1).
@@ 2.2 Inference Stage
Different queries require different perspectives to answer effectively. For instance, queries about global understanding, like “What is the overall theme of this presentation?” requires a broad overview from a macroscopic perspective and activates only the global agent. In contrast, a fact-based query, like “What is the revenue on slide 3?” requires detailed, slide-specific information and requires both the page and element agents. Leveraging too many agents may increase computation or introduce noise. Thus, the agent orchestrator first attempts to classify each query $q$ into one of 4 predefined categories, such as global understanding, fact-based direct queries, multi-hop reasoning, and layout & visual relationships. Each type corresponds to a question-answering strategy and a targeted set of agents (Table 1). For instance, global understanding queries activate only the global agent, as they require a broad overview, while fact-based queries trigger both the page and element agents to retrieve detailed, slide-specific information. If none of the predefined categories apply, the agent defaults to the “unknown” category, which activates all agents.
@@ 3 Experiments
Gemini-2.0-Flash 75.0 71.3 79.8 50.4 67.6 41.6 70.8 70.6 77.8
Gemini-2.5-Flash 83.8 78.3 91.8 51.1 71.4 41.2 76.2 75.8 100.0
Gemini-2.5-Flash-Lite 71.2 60.8 87.0 47.3 58.1 41.9 57.0 56.6 68.3
Claude-4.1-Opus 78.4 74.3 82.3 61.0 81.4 52.3 56.5 54.8 73.3
Claude-3.5-Haiku 62.5 68.3 54.6 52.5 80.2 39.5 48.5 49.5 29.6
@@
Multimodal RAG (Type 2) and Agentic Methods (Type 3) with GPT-4o as base models.
@@
Table 2: Performance comparison of proprietary models on SlideVQA, TechSlides, and FinSlides. All baseline methods use GPT-4o for answer generation. Improvements over the base model GPT-4o is shown in green. The best and second best performance are highlighted in bold and underlined. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same base model, and even outperforms stronger raw LLMs (e.g. Gemini-2.5). late
@@
We benchmark SlideAgent against 3 types of baselines: 1) Multimodal LLMs: 15 MLLMs from 8 model families, including proprietary models (GPT-4o (OpenAI, 2025), Gemini (Anil et al., 2023), Claude (Anthropic, 2025)) and open-source models (Llama-3.2 (Grattafiori et al., 2024), InternVL3 (Chen et al., 2024), Phi-3 (Abdin et al., 2024), Qwen2.5-VL (Bai et al., 2025)); 2) Multimodal RAG Methods: VisRAG Yu et al. (2025), VDocRAG Tanaka et al. (2025), and COLPALI Faysse et al. (2024); 3) Multi-agent Systems: ViDoRAG Wang et al. (2025a) and MDocAgent (Han et al., 2025). We evaluate SlideAgent using two backbone MLLMs, including both proprietary models (GPT-4o (OpenAI, 2025)) and open-source models (InternVL3-8B (Chen et al., 2024)). These models are chosen for their widespread use in state-of-the-art QA systems Yu et al. (2025); Jin et al. (2025); Cho et al. (2024). The parameter sizes, knowledge cutoff dates, and release dates are in Appendix Table 9. We use the text-based retriever SFR (Meng et al., 2024) due to its strong efficiency-performance Cheng et al. (2024); Jin et al. (2025). For models restricted to single-image input (e.g. LLaVA-v1.5 (Liu et al., 2023)), we concatenate the top 3 retrieved images into one following previous work Yu et al. (2025). In settings where ground-truth pages are available, we provide them as input to all models. Otherwise, each model receives as many retrieved pages up to its input capacity.
@@
Multimodal LLMs
Gemini-2.0-Flash 72.3 66.1 81.3
Gemini-2.5-Flash 86.9 81.1 95.4
Gemini-2.5-Flash Lite 71.7 62.6 87.0
Claude-4.1-Opus 36.3 35.4 38.9
Claude-3.5-Haiku 31.5 30.8 37.0
@@
Table 3: Performance comparison among proprietary models, such as Gemini, Claude, GPT-4o, ViDoRAG, and SlideAgent on InfoVQA Mathew et al. (2022).
@@ 3.1 End-to-End Performance
Tables 2/3/4 and Appendix Table 5 demonstrate the end-to-end performance of SlideAgent across proprietary and open-source models.
@@
Llama-3.2-11B-Vision-Instruct 42.9 43.3 42.3 41.4 52.5 36.2 23.3 23.2 26.2
Phi-3-vision-128k-instruct 72.3 61.8 90.6 59.4 60.0 59.1 48.8 48.5 64.3
Qwen2.5-VL-7B-Instruct 79.5 70.5 94.3 59.3 52.5 65.9 53.6 52.5 85.7
Qwen2.5-VL-32B-Instruct 79.2 71.1 92.2 67.5 87.5 60.6 57.4 56.6 87.5
llava-1.5-7b-hf 36.8 22.4 79.0 23.3 12.5 38.7 10.7 10.1 16.6
llava-1.5-13b-hf 44.9 25.1 81.8 28.1 17.5 45.0 20.6 16.5 36.7
llava-v1.6-mistral-7b-hf 50.9 37.3 82.6 34.4 37.5 32.2 12.2 12.1 17.8
llava-v1.6-vicuna-13b-hf 16.7 10.2 81.5 45.2 40.0 49.1 32.0 31.3 64.3
InternVL3-8B 63.0 56.5 74.1 55.4 57.5 54.4 49.8 49.5 64.3
Multimodal RAG (Type 2) and Agentic Method (Type 3) with InternVL3-8B as base models.
@@
Table 4: Performance (correctness %) of various models on SlideVQA, TechSlides, and FinSlides dataset. SlideAgent outperforms all baseline methods (Type 2/3) sharing the same LLM backbone.
@@
For both proprietary and open-source models, SlideAgent consistently outperforms all baseline methods (Type 2/3) and the base models across all metrics. For proprietary models (Table 2), SlideAgent improves overall accuracy by +7.9+7.9%, numeric reasoning by +8.3+8.3%, and lexical overlap by +6.5+6.5% on SlideVQA. Similar trends are observed for TechSlides (+7.5+7.5% overall; +12.3+12.3% F1) and FinSlides (+5.5+5.5% overall; +17.5+17.5% F1), indicating robust performance across diverse domains. For open-source models, the advantage remains pronounced: SlideAgent outperforms the original MLLM by +9.8+9.8% overall (+11.7+11.7% numeric), showing that our hierarchical, multi-agent design generalizes well across MLLMs.
Figure 3: Absolute performance gains over MDocAgent Han et al. (2025) when using GPT-4o and InternVL3-8B as the backbone models on SlideVQA Tanaka et al. (2023). SlideAgent yields significant overall improvements across both backbones.
@@
SlideAgent consistently achieves the best or second-best performance among proprietary models (Table 2 and Appendix Table 11). Notably, although the base model slightly lags behind stronger MLLMs such as Gemini-2.5-Flash in raw capability (e.g. 83.8 for Gemini-2.5-Flash vs. 77.077.0% for GPT-4o in Table 2), SlideAgent’s structured reasoning pipeline fills this gap (84.984.9% overall for SlideAgent). For open-source models, SlideAgent achieves better performance across all models except for the Qwen2.5-VL family. While strong base models such as Gemini-2.5 and Qwen2.5-VL already exhibit advanced multimodal comprehension, SlideAgent is model-agnostic in nature and can be directly applied to these models to further enhance performance.
@@
Figure 4 presents the performance breakdown across query types, as defined in Table 1. SlideAgent improves question-answering across diverse query types, especially in multi-hop reasoning and visual/layout questions. The greatest improvement occurs on case 3 (multi-hop reasoning), with a 9.89.8% improvement (67.467.4% to 77.277.2%). This means explicitly guiding the model’s reasoning using generated knowledge $(\mathcal{K}_{g},\mathcal{K}_{p},\mathcal{K}_{e})$ significantly improves reasoning capabilities. A notable 7.77.7% improvement (66.766.7% $\rightarrow 74.4$%) is also observed in visual/layout reasoning, demonstrating the benefit of fine-grained element-level reasoning and retrieval that multimodal LLMs hardly achieve. Both SlideAgent and its counterpart perform well on Case 2 (Fact-based Direct Queries), with only a modest 2.12.1% improvement, reflecting the model’s ability to handle page-level reasoning with little space for further enhancement.